he followi Chapter 3 Homework Check my work mode: This shows what is correct or incorrect for the work you h Required information The following information applies to the questions displayed below. The following information pertains to Trenton Glass Works for the year just ended Part 2 of 3 Budgeted direct-labor cost 70,000 hours (practical capacity) at $16 per hour Actual direct-labor cost:80,000 hours at $17.50 per hour Budgeted manufacturing overhead: $997500 Actual selling and administrative expenses: 438,000 0.76 points Act ual manufacturing overhead: Depreciation Property taxes Indirect labor Supervisory salaries Utilities Insurance Rental of space Indirect material (see data below) $234,000 21,000 83,000 202,000 59,000 31,000 301,000 78,000 Indirect material: Beginning inventory, January 1 Purchases during the year Ending inventory, December 31 47,000 94,000 63,000 2. Calculate the overapplied or underapplied overhead for the year. (Round your intermediate calc Answer is complete but not entirely correct. Overapplied overhead 391,000
Answer #1

Solution: 131,000 (Over-applied)

Working:

Actual manufacturing overhead

Cost ($)

Depreciation

$234,000

Property Taxes

$21,000

Indirect Labor

$83,000

Supervisory salaries

$202,000

Utilities

$59,000

Insurance

$31,000

Rental of Space

$301,000

Indirect Material

$78,000

Total Actual Manufacturing Overhead

$1,009,000

Predetermined Overhead Rate = Budgeted Manufacturing overhead cost / Budgeted amount of cost driver (or activity based) = 997,500 / 70,000 = 14.25

Applied overhead = Actual labor hours worked x predetermined overhead rate = 80,000 * 14.25 = 1,140,000

Over/Under Applied = Applied Overhead – Actual Manufacturing Overhead = 1,140,000 - 1,009,000 = 131,000 (Over-applied)